OpenStudy (leonard):

What is the solution of the system? Use substitution. 3x + 2y = 7 y = –3x + 11 A. (6, –3) B. (6, –7) C. D. (5, –4)

OpenStudy (legomyego180):

Do you know how to start this

563blackghost (563blackghost):

We know that y equals \(\large{-3x+11}\) for the second equation...since we know of y we can substitute this equation where y is in the first equation soo... \(\Huge{3x+2(-3x+11)=7}\) In which you would simplify to find x...

OpenStudy (triciaal):

@Er.Mohd.AMIR the "asker" is offline. Do you have a question about this?

OpenStudy (triciaal):

what?

OpenStudy (er.mohd.amir):

i know how to solve it i have no q about this one

OpenStudy (triciaal):

ok bye

OpenStudy (aaronandyson):

3x + 2y = 7(eqn1) y = –3x + 11(eqn2) 3x + 2(-3x + 11) = 7 3x -6x + 22 = 7 -3x +22 = 7 -3x = 7-22 -3x = -15 3x = 15 x = 15/3 x = 5 Now ,we either place x = 5 in eqn1 or eqn2 I'll put in eqn1 3x + 2y = 7 3(5) + 2y = 7 15 + 2y = 7 2y = 7-15 2y = -8 y =-8/2 y = -4 Therefore the solution is (5,-4)